OpenStudy (anonymous):

What is C would mean that the system of equation 2x-3y=1 and cx-3y=2 would not have any solutions

OpenStudy (anonymous):

if both coefficients were the same the lines would be parallel

OpenStudy (anonymous):

I don't know what that means

OpenStudy (anonymous):

i mean if you have \[2x-3y=1\] and \[2x-3y=2\] then the lines are parallel

OpenStudy (anonymous):

Then had did you know the c was 2

OpenStudy (anonymous):

i am not sure what you are asking

OpenStudy (anonymous):

Well the second equation had cx but then you substituted it as a 2

OpenStudy (anonymous):

Had did you know that

OpenStudy (texaschic101):

Let me try to explain....when lines have the same slope but different y intercepts, then they are parallel lines and will have no solutions. 2x - 3y = 1 (put this in y = mx + b form where the number in the m position is the slope and the number in the b position is the y intercept) -3y = -2x + 1 -- divide by -3 y = (-2/-3) x + 1/3 y = 2/3x + 1/3 so our slope here is 2/3 and our y intercept is 1/3 cx - 3y = 2 -3y = -cx + 2 -- divide by -3 y = (-c/-3) - 2/3 y = c/3 - 2/3 now we have to have the same slope in order for them to be parallel. So c will have to be 2 because the slope of the first line was 2/3. And since the y intercepts are both different....that makes the lines parallel, thus, they have no solution

OpenStudy (anonymous):

The answer would be 2 even if there is no solution

OpenStudy (texaschic101):

if c = 2, then the lines are parallel and have no solution. However, if c would have been a different number, such as 4, then the lines would not be parallel because parallel lines have the same slope, thus, there would be 1 solution. Learn this...it will help if the slopes and the y intercepts are the same, then the lines coincide (they lay on top of each other) and they will have infinite solutions. if the slopes are the same, but the y intercepts are different, then the lines are parallel and they never intersect, thus, they have no solution. if the slopes and the y intercepts are different, then the lines will intersect, thus having 1 solution.

OpenStudy (anonymous):

Oh ok now I get it